OpenStudy (raffle_snaffle):

Given points A(1,-5,2), B(3,5,0), and C(-1,2,-7) in 3-dimension: a. Find a unit vector orthogonal to the plane containing A, B, and C. b. Find the equation, in rectangular coordinates, of the plane through A, B, and C.

OpenStudy (raffle_snaffle):

So part (a) I got n=<104, 14, 34>

OpenStudy (psymon):

Whoa, whoa, whoa, how is a "unit" vector, going to have such massive numbers? :P

OpenStudy (raffle_snaffle):

and part (b) I got 104(x-1)+14(y+5)+34(z-2)=0

OpenStudy (raffle_snaffle):

well that is my question. I don't think i did part a right.Lol

OpenStudy (psymon):

Well, recall that a unit vector is found by dividing your vector by its magnitude.

OpenStudy (raffle_snaffle):

... ugh

OpenStudy (raffle_snaffle):

i remember you tellling me that a few days ago.

OpenStudy (psymon):

Yep. Well let's just make sure you did the firstpart correctly. So to find a vector orthogonal, what did you do?

OpenStudy (raffle_snaffle):

v=<2sqrt2/3, 7sqrt2/78, 17sqrt2/8>

OpenStudy (psymon):

Lol, I guess I should just check it instead of asking what you did xDD

OpenStudy (raffle_snaffle):

you want me to find the normal vector?

OpenStudy (raffle_snaffle):

I would appreciate that because I don't want to show all my work. Lol

OpenStudy (psymon):

No worries. I was only asking for a verbal explanation is all, dont expect ya to type out all the work xD

OpenStudy (psymon):

Ouch, big numbers.

OpenStudy (raffle_snaffle):

so?

OpenStudy (psymon):

Your first answer that had large numbers, it looks like you did the same thing I did, but we had some different signs that resulted in different answers. Just checking my cross-product work.

OpenStudy (raffle_snaffle):

ab=<2,10,2> or <2,10,-2>

OpenStudy (raffle_snaffle):

at first I had <2,10,-2> but then i made a little error, suppose to be <2,10,2>

OpenStudy (psymon):

No, the very first thing you posted that had <104, 14, 34> Looks like you did some mistakes with signs.

OpenStudy (psymon):

And it should be -2 based on what you typed your points to be.

OpenStudy (raffle_snaffle):

but it's 0-(-2)=2

OpenStudy (psymon):

Then you typed your point up top wrong.

OpenStudy (raffle_snaffle):

so it's suppose to be -2?

OpenStudy (psymon):

You typed your point as <1, -5, 2>

OpenStudy (psymon):

So that means it would be (0-2) = -2.

OpenStudy (raffle_snaffle):

when i am looking for AB you are suppose to take b-a to obtain the vector

OpenStudy (psymon):

Correct.

OpenStudy (raffle_snaffle):

my given points are right. I just looked them over.

OpenStudy (psymon):

Well, if the point <1, -5, 2> is correct, then it'd be 0 - 2

OpenStudy (raffle_snaffle):

okay let me fix it..

OpenStudy (psymon):

So the point should be <1, -5,-2> ?

OpenStudy (raffle_snaffle):

n=<-76, 22, 34>

OpenStudy (psymon):

Bleh, Ill just assume the point is <1. -5. -2> x_x *reworks*

OpenStudy (raffle_snaffle):

no it's (1,-5, 2)

OpenStudy (psymon):

lol x_x Nvm then. Okay, hang on o.o

OpenStudy (psymon):

Then yes, your normal vector is correct, now you just need to divide it by its magnitude.

OpenStudy (raffle_snaffle):

kk

OpenStudy (raffle_snaffle):

<-19sqrt206/309, 11sqrt206/618, 17sqrt206/618>

OpenStudy (psymon):

Alright, awesome.

OpenStudy (psymon):

So now we just need the equation of the plane :P

OpenStudy (raffle_snaffle):

sorry i lost connection..

OpenStudy (raffle_snaffle):

okay equation of the plane is -76(x-1)+22(y+5)+34(z-2)=0

OpenStudy (psymon):

Seems good to me :3

OpenStudy (raffle_snaffle):

it doesn't check though

OpenStudy (psymon):

Oh?

OpenStudy (raffle_snaffle):

hold on

OpenStudy (raffle_snaffle):

nvm it does thanks man!

OpenStudy (psymon):

Lol, okay, awesome xD Yeah, you had the right idea, just had to remember how to get the unit vector :P

OpenStudy (raffle_snaffle):

ive got one more question

OpenStudy (psymon):

kk

OpenStudy (raffle_snaffle):

#17 find the vector equation of the line containing the points (3,-1,1) and (4,0,2). Determine where, if at all, the line intersects the xy, yz, and xz-planes. #18 Given points A(1,-5,2), B(3,5,0), and C(-1,2,-7) in 3-dimension: a. Find a unit vector orthogonal to the plane containing A, B, and C. b. Find the equation, in rectangular coordinates, of the plane through A, B, and C. Main question; Determine where the line in #17 intersects the plane in #18; if they do not intersect, state this.

OpenStudy (raffle_snaffle):

vector equation is v=<3+t, -1-t, 1+t>

OpenStudy (raffle_snaffle):

i already answered #17 and #18

OpenStudy (psymon):

So only the main question is left?

OpenStudy (raffle_snaffle):

yes

OpenStudy (psymon):

I havent had a question that asked those specific things, so not exactly sure if Id know how to do it xD But then again I dont know if I do or dont yet. So whats the plane equation since you already did it?

OpenStudy (raffle_snaffle):

plane equation is -76x+22y+34z=-118

OpenStudy (psymon):

Oh, what we just did./

OpenStudy (raffle_snaffle):

yeahxD

OpenStudy (raffle_snaffle):

btw that last question you asked me a few days ago, I can probably help you. I had a problem just like it in my text book.

OpenStudy (psymon):

Yeah, and I dont think ive had a problem like this in mine xDD Ill have to see if I can figure out how to find that intersection.

OpenStudy (raffle_snaffle):

well the equation of the plane is wrong... I was looking at the key my instructor posted

OpenStudy (psymon):

We still need to know how to check the intersection regardless xD

OpenStudy (raffle_snaffle):

it was suppose to be those larger numbers... lol

OpenStudy (raffle_snaffle):

so the-2 was suppose to be a +2

OpenStudy (psymon):

Meh, I dont care at this point, haha. Well if you have the correct plane then we should see how we get the intersection.

OpenStudy (raffle_snaffle):

correct plane 104x-14y+34z=242

OpenStudy (psymon):

Im not sure how we do a system with a line equation and a plane equation xD

OpenStudy (raffle_snaffle):

okay lets push this question aside. I've got one more question

OpenStudy (psymon):

Alrighty. Someone else would know of course.

OpenStudy (raffle_snaffle):

Explain how to find the displacement vector between two points in 3-space, and then find the unit vector in the same direction. Give an example.

OpenStudy (psymon):

Bleh, im getting confused on my own work, lol.

OpenStudy (raffle_snaffle):

no worries. This material can be quite confusing.

OpenStudy (psymon):

I cant tell when I need to use cross-product and when i don't. I have some problems where it wants parallel planes and lines and uses cross-product and I have some where it wants perpendicular planes and lines and still uses cross-product.

OpenStudy (raffle_snaffle):

callisto scroll down to the last question

OpenStudy (psymon):

Bleh, I know how to do the work, just dont know when I need cross-product and when I don't -_-

OpenStudy (raffle_snaffle):

so lets use the points in the previous question

OpenStudy (psymon):

And find where it intersects the plane, right?

OpenStudy (raffle_snaffle):

no no not that question

OpenStudy (raffle_snaffle):

This one: Explain how to find the displacement vector between two points in 3-space, and then find the unit vector in the same direction. Give an example.

OpenStudy (psymon):

By displacement vector is that the same as component vector?

OpenStudy (raffle_snaffle):

i guess

OpenStudy (raffle_snaffle):

displacement is final minus intital

OpenStudy (psymon):

Right. Final of the corresponding directions. i final minus i initial, etc.

OpenStudy (raffle_snaffle):

so if we chose two points in space A(1,-5,2) and B(3,5,0)...

OpenStudy (raffle_snaffle):

we get AB=<2,10,-2>

OpenStudy (psymon):

Pretty easy for this question. (3-1, 5-(-5), 0-2) = (2, 10, -2) sqrt(4 + 100 + 4) = sqrt(108) = 6sqrt(3) <1/3sqrt(3), 5/3sqrt(3), -1/3sqrt(3)>

OpenStudy (raffle_snaffle):

so the unit vector is taking the vector we obtained from the x product and then dividing it by the magnitude of the vector we found?

OpenStudy (psymon):

It only wanted the unit vector in the same direction as the component vector, so we don't need a cross product for that.

OpenStudy (raffle_snaffle):

ah okay

OpenStudy (psymon):

So we just took the component vector and divided it by the magnitude.

OpenStudy (raffle_snaffle):

callisto i can't reply to your email..

OpenStudy (raffle_snaffle):